OpenStudy (sugarmochi):

Find the parabola that passes through the set of points: a (-2,-3) b (1,0) c (3,32)

OpenStudy (steve816):

Looks like you need to solve system of equations with 3 variables! Yayy fun :)

OpenStudy (steve816):

Note that the general form of the parabola is \[\large y=ax^2+bx+c\] First step: you want to plug in your x coordinates into x in the equation and the y coordinates into y in the equation. You should have 3 different equations.

OpenStudy (sugarmochi):

so the equations would be -3=-4a+b+c 0=a+b+c 32=9a+3b+c

OpenStudy (steve816):

The first equation should be -3=4a-2b+c

OpenStudy (sugarmochi):

ahh my bad

OpenStudy (steve816):

Now, I don't know what the best method is to solve that 3 variable system of equations, but I would try substitution method.

OpenStudy (sugarmochi):

yeah thankyou ! I got this now, i was originally confused cos it wasnt in the table format that i was used to but this cleared it up. THanks !

OpenStudy (steve816):

Hey, no problem! Good luck on solving the rest :)